OpenStudy (anonymous):

if y=sin(m sin^-1x),then prove,(1-x^2)y2-xy1+m^2y=0 ???????

OpenStudy (anonymous):

y1=d/dx(sin(m sin^-1x) =cos(m sin^-1x) multiple m/root 1-x^2 =cos^2(m sin^-1x) multiple m^2/(1-x)^2 y1^2(1-x)^2=m^2cos^2(m sin ^-1x) =m^2[1-sin^2(m sin^-1x] m^2[1-y^2] or(1-x)^22y1 y2-2xy1^2=-2m^2y y1 (1-x^2)y2-xy1+m^2y=0

OpenStudy (anonymous):

thanks everyone 4 helping me!!!!!!!!